# Fast Matrix Multiplications for Lookup Table-Quantized LLMs

### v1

@@ 1 Introduction
Realizing practical speed-ups with weight-only quantization requires custom mixed-type matrix-matrix multiply (matmul) kernels which must (1) move a layer’s quantized weights from GPU off-chip DRAM to on-chip SRAM, (2) dequantize the weights to floating-point (FP) format (on chip), (3) perform the FP matmul, and (4) write the results back to DRAM. Existing kernels such as bitsandbytes Dettmers et al. (2023), Marlin Frantar and Alistarh (2024), and BitBLAS Wang et al. (2024) demonstrate that this strategy can result in significant matmul speed-ups, e.g. up to four times faster when going from W16A16 to W4A16. However, these kernels are typically specialized to 4-bit quantization, and while some kernels support non-uniform, lookup table (LUT) quantization, they are generally slower than the uniform counterparts. Given the recent promising results with odd-bit Shao et al. (2023); Ma et al. (2024b, a) and non-uniform Guo et al. (2024); Kim et al. (2023) quantization methods, there is thus a need to develop flexible kernels that can support mixed-type matmuls with a wider range of settings.
@@
FLUTE addresses these challenges through a combination of (1) offline weight restructuring, (2) a shared-memory lookup table for efficient dequantization, and (3) Stream-K partitioning for optimized workload distribution. We compare FLUTE against existing kernels on standard LLM mixed-precision matmul settings where weights are quantized to 4 bits in groups of 128, and find that it outperforms existing non-uniform quantization kernels, and even matches the simpler uniform-quantization kernels in some cases. As an application of FLUTE, we experiment with quantizing LLaMA3—which has been found to be difficult to quantize Huang et al. (2024)—using a variant of normal float (NF) quantization Dettmers et al. (2023) which learns the quantization parameters based on calibration data. We find that we can achieve an 1.5 to 2 times increase in end-to-end throughput when integrated with frameworks such as vLLM Kwon et al. (2023).
@@ 2.2 LLM Deployment Characteristics
Depending on the context, inference can be bottlenecked by compute throughput or memory bandwidth. For LLMs, training, large-prefill, and large-batch inference enjoy high arithmetic intensity as the sizes of matrices involved in the matmuls are large enough to saturate compute. Small-batch, token-by-token inference on the other hand involves narrower matmuls due to the smaller batch dimension, resulting in low arithmetic intensity. Reducing the amount of memory operations in this case can thus enable practical speed-ups, even if the number of FLOPs remains the same (or is even slightly increased). This has led to much recent work on customized kernels which move the weights from main memory to on-chip SRAM while keeping them quantized/sparse Dettmers et al. (2023); Kim et al. (2023); Frantar and Alistarh (2024); Wang et al. (2024); Xia et al. (2024a), and then performing the actual matmuls in higher precision after dequantizing to FP on chip. Marlin implements this strategy for 4-bit uniform quantization and reports significant (up to 4$\times$) matmul speed-ups even in moderate batch (16-32) settings. bitsandbytes Dettmers et al. (2023) and BitBLAS Wang et al. (2024) extend this to LUT-quantized LLMs, but do not allow for 3 bit-quantized weights. Moreover, existing LUT-quantization kernels generally underperform uniform-quantization kernels.
@@ 3.1 Offline Matrix Restructuring
Modern GPUs feature specialized primitives (Tensor Cores)—distinct from general-purpose vector ALUs—which can substantially accelerate dense matrix multiplications. For example, A100’s FP16 tensor core matmuls are 16$\times$ faster than FP32 vector matmuls. However, this acceleration comes at the expense of generality and programmability. Tensor Core MMA (matrix-multiply-accumulate) operations require the input matrices to adhere to specific layout specifications within the registers of 32 threads. The fused kernel needs to first load fragments of the quantized weight into registers, dequantize the matrix, and then perform the MMA operation between the input fragments and dequantized matrix fragments. This necessitates that the post-dequantization matrix layout meets the required specifications. While runtime data reordering is one approach, it introduces a substantial number of operations. Instead, we leverage the fact that ${\mathbf{Q}}$ (the quantized weights) are static during inference, allowing for offline weight reordering such that after dequantization, the weights are already laid out exactly in the expected format Frantar and Alistarh (2024); Xia et al. (2024b); Lin et al. (2024).
@@ 4.1 Kernel Benchmarks
There are many methods for LUT quantization; we follow the popular NormalFloat LUT quantization scheme Dettmers et al. (2023), where a tensor-level table $\mathbf{T}$ is modified to be a group-level table via a scaling factor $s_{g}$ for each group $g$, resulting in a group-level table $\mathbf{T}_{g}=[s_{g}\cdot v_{0},\dots,s_{g}\cdot v_{2^{b}-1}]$. Here $s_{g}\in\mathbb{R}^{+}$ is a scalar that varies per group. This approach requires maintaining a tensor-level lookup table $\mathbf{T}$ and group-level scalars $\{s_{g}\}_{g=1}^{d^{2}/g}$, and thus incur almost the same memory overhead as uniform quantization (which also requires maintaining the group-level scalars). While we primarily focus on LUT kernels, for completeness we also compare against high-performance kernels specialized for uniformly quantized weights (BitBLAS [^3] https://github.com/microsoft/BitBLAS and Marlin [^4] https://github.com/IST-DASLab/marlin). These kernels do not require dynamic indexing into a lookup table and can perform dequantization in registers using highly tuned PTX assembly instructions that are not applicable to LUT-based dequantization.
@@ 5 Discussion and Conclusion
Early work on LLM quantization generally worked with uniform quantization methods Frantar et al. (2022); Dettmers et al. (2022); Xiao et al. (2022). More recent work has shown the benefits of LUT-quantization, both from PTQ Kim et al. (2023) and finetuning Dettmers et al. (2023) perspectives. Insofar as lookup tables can represent flexible quantization functions, our hope is that FLUTE can enable researchers and practitioners to explore new quantization algorithms that can learn better lookup tables Yamamoto (2021); Cardinaux et al. (2020); Wang et al. (2022). On a similar note, recent work has found that codebook-based quantization schemes—which generalize lookup tables to vector-valued values—can enable even lower-bit (e.g., 2-bit) LLM quantization without significant performance degradations Tseng et al. (2024); Egiazarian et al. (2024). We anticipate that ideas from this work can aid in developing kernels for such methods.
@@ Acknowledgements
We thank Yijie Bei and Dmytro Ivchenko for helpful discussion. HG was supported by a Microsoft PhD Fellowship. EX acknowledges the support of NGA HM04762010002, NSF IIS1955532, NSF CNS2008248, NIGMS R01GM140467, NSF IIS2123952, NSF DMS2027737, NSF BCS2040381, NSF DMS2112273, NSF IIS2311990, Semiconductor Research Corporation (SRC) AIHW award 2024AH3210, and DARPA ECOLE HR00112390063. This study was additionally supported by funds from MIT-IBM Watson AI and the MLA@CSAIL initiative.

### v3

@@ 1 Introduction
Realizing practical speed-ups with weight-only quantization requires custom mixed-type matrix-matrix multiply (matmul) kernels which must (1) move a layer’s quantized weights from GPU off-chip DRAM to on-chip SRAM, (2) dequantize the weights to floating-point (FP) format (on chip), (3) perform the FP matmul, and (4) write the results back to DRAM. Existing kernels such as bitsandbytes Dettmers et al. (2023), Marlin Frantar et al. (2024), and BitBLAS Wang et al. (2024) demonstrate that this strategy can result in significant matmul speed-ups, e.g. up to four times faster when going from W16A16 to W4A16. However, these kernels are typically specialized to 4-bit quantization, and while some kernels support non-uniform, lookup table (LUT) quantization, they are generally slower than the uniform counterparts. Given the recent promising results with odd-bit Shao et al. (2023); Ma et al. (2024b, a) and non-uniform Guo et al. (2024); Kim et al. (2023) quantization methods, there is thus a need to develop flexible kernels that can support mixed-type matmuls with a wider range of settings.
@@
FLUTE addresses these challenges through a combination of (1) offline weight restructuring, (2) a shared-memory lookup table for efficient dequantization, and (3) Stream-K partitioning for optimized workload distribution. We compare FLUTE against existing kernels on standard LLM mixed-precision matmul settings where weights are quantized to 4 bits in groups of 128, and find that it outperforms existing non-uniform quantization kernels, and even matches the simpler uniform-quantization kernels in some cases. As an application of FLUTE, we experiment with quantizing LLaMA3—which has been found to be difficult to quantize Huang et al. (2024)—using a variant of normal float (NF) quantization Dettmers et al. (2023) which learns the quantization parameters based on calibration data. We find that we can achieve a 1.5 to 2 times increase in end-to-end throughput when integrated with frameworks such as vLLM Kwon et al. (2023).
@@ 2.2 LLM Deployment Characteristics
Depending on the context, inference can be bottlenecked by compute throughput or memory bandwidth. For LLMs, training, large-prefill, and large-batch inference enjoy high arithmetic intensity as the sizes of matrices involved in the matmuls are large enough to saturate compute. Small-batch, token-by-token inference on the other hand involves narrower matmuls due to the smaller batch dimension, resulting in low arithmetic intensity. Reducing the amount of memory operations in this case can thus enable practical speed-ups, even if the number of FLOPs remains the same (or is even slightly increased). This has led to much recent work on customized kernels which move the weights from main memory to on-chip SRAM while keeping them quantized/sparse Dettmers et al. (2023); Kim et al. (2023); Frantar et al. (2024); Wang et al. (2024); Xia et al. (2024a), and then performing the actual matmuls in higher precision after dequantizing to FP on chip. Marlin implements this strategy for 4-bit uniform quantization and reports significant (up to 4$\times$) matmul speed-ups even in moderate batch (16-32) settings. bitsandbytes Dettmers et al. (2023) and BitBLAS Wang et al. (2024) extend this to LUT-quantized LLMs, but do not allow for 3 bit-quantized weights. Moreover, existing LUT-quantization kernels generally underperform uniform-quantization kernels.
@@ 3.1 Offline Matrix Restructuring
Modern GPUs feature specialized primitives (Tensor Cores)—distinct from general-purpose vector ALUs—which can substantially accelerate dense matrix multiplications. For example, A100’s FP16 tensor core matmuls are 16$\times$ faster than FP32 vector matmuls. However, this acceleration comes at the expense of generality and programmability. Tensor Core MMA (matrix-multiply-accumulate) operations require the input matrices to adhere to specific layout specifications within the registers of 32 threads. The fused kernel needs to first load fragments of the quantized weight into registers, dequantize the matrix, and then perform the MMA operation between the input fragments and dequantized matrix fragments. This necessitates that the post-dequantization matrix layout meets the required specifications. While runtime data reordering is one approach, it introduces a substantial number of operations. Instead, we leverage the fact that ${\mathbf{Q}}$ (the quantized weights) are static during inference, allowing for offline weight reordering such that after dequantization, the weights are already laid out exactly in the expected format Frantar et al. (2024); Xia et al. (2024b); Lin et al. (2024).
@@ 4.1 Kernel Benchmarks
There are many methods for LUT quantization; we follow the popular NormalFloat LUT quantization scheme Dettmers et al. (2023), where a tensor-level table $\mathbf{T}$ is modified to be a group-level table via a scaling factor $s_{g}$ for each group $g$, resulting in a group-level table $\mathbf{T}_{g}=[s_{g}\cdot v_{0},\dots,s_{g}\cdot v_{2^{b}-1}]$. Here $s_{g}\in\mathbb{R}^{+}$ is a scalar that varies per group. This approach requires maintaining a tensor-level lookup table $\mathbf{T}$ and group-level scalars $\{s_{g}\}_{g=1}^{d^{2}/g}$, and thus incur almost the same memory overhead as uniform quantization (which also requires maintaining the group-level scalars). While we primarily focus on LUT kernels, for completeness we also compare against high-performance kernels specialized for uniformly quantized weights (BitBLAS [^3] https://github.com/microsoft/BitBLAS and Marlin [^4] https://github.com/IST-DASLab/marlin Frantar et al. (2024)). These kernels do not require dynamic indexing into a lookup table and can perform dequantization in registers using highly tuned PTX assembly instructions that are not applicable to LUT-based dequantization.
@@ 4.2 End-to-End LLM Benchmarks
Figure 6: End-to-end vLLM latency benchmark for processing a single batch of requests using LLaMA-3.1 (405B, W4G64).
@@ Results.
To demonstrate the scalability of our approach, we evaluated FLUTE on the LLaMA-3.1 (405B) model Dubey et al. (2024), shown in Fig. 6. It is worth noting that, without quantization, the 405B model’s parameters alone would require multiple GPU nodes. However, with FLUTE, we were able to perform inference on a single node, highlighting the efficiency and scalability of our solution.
@@ 5 Discussion and Conclusion
Early work on LLM quantization generally worked with uniform quantization methods Frantar et al. (2022); Dettmers et al. (2022); Xiao et al. (2022). More recent work has shown the benefits of LUT-quantization, both from PTQ Kim et al. (2023) and finetuning Dettmers et al. (2023) perspectives. Insofar as lookup tables can represent flexible quantization functions, our hope is that FLUTE can enable researchers and practitioners to explore new quantization algorithms that can learn better lookup tables Yamamoto (2021); Cardinaux et al. (2020); Wang et al. (2022). For example, recent work has found that codebook-based quantization schemes—which generalize lookup tables to vector-valued values—can enable even lower-bit (e.g., 2-bit) LLM quantization without significant performance degradations Tseng et al. (2024); Egiazarian et al. (2024). We anticipate that ideas from this work can aid in developing kernels for such methods.
@@ Acknowledgements
We thank Yijie Bei and Dmytro Ivchenko for helpful discussion. HG was supported by a Microsoft PhD Fellowship. EX acknowledges the support of NGA HM04762010002, NSF IIS1955532, NSF CNS2008248, NIGMS R01GM140467, NSF IIS2123952, NSF DMS2027737, NSF BCS2040381, NSF DMS2112273, NSF IIS2311990, Semiconductor Research Corporation (SRC) AIHW award 2024AH3210, and DARPA ECOLE HR00112390063. This study was additionally supported by MIT-IBM Watson AI Lab and the MLA@CSAIL initiative.